# Zhening Huang1       Xiaoyang Wu2       Fangcheng Zhong1 Hengshuang Zhao2       Matthias Nießner3       Joan Lasenby1

### v1

@@ Abstract
We propose LiteReality, a novel pipeline that converts RGB-D scans of indoor environments into compact, realistic, and interactive 3D virtual replicas. LiteReality not only reconstructs scenes that visually resemble reality but also supports key features essential for graphics pipelines—such as object individuality, articulation, high-quality physically based rendering materials, and physically based interaction. At its core, LiteReality first performs scene understanding and parses the results into a coherent 3D layout and objects, with the help of a structured scene graph. It then reconstructs the scene by retrieving the most visually similar 3D artist-crafted models from a curated asset database. Later, the Material Painting module enhances the realism of retrieved objects by recovering high-quality, spatially varying materials. Finally, the reconstructed scene is integrated into a simulation engine with basic physical properties applied to enable interactive behavior. The resulting scenes are compact, editable, and fully compatible with standard graphics pipelines, making them suitable for applications in AR/VR, gaming, robotics, and digital twins. In addition, LiteReality introduces a training-free object retrieval module that achieves state-of-the-art similarity performance, as benchmarked on the Scan2CAD dataset, along with a robust Material Painting module capable of transferring appearances from images of any style to 3D assets—even in the presence of severe misalignment, occlusion, and poor lighting. We demonstrate the effectiveness of LiteReality on both real-life scans and public datasets.
Figure 1: From RGB-D Scan to Graphics-Ready 3D Scene. LiteReality reconstructs compact, realistic 3D environments from real-world RGB-D scans, featuring near-photorealistic appearance, high-quality articulated geometry, and physically based rendering (PBR) materials—ready for high-fidelity rendering and interactive simulation.
@@ 1 Introduction
Creating digital replicas of real-world environments remains a central challenge in machine perception, computer vision, and graphics. Although differentiable rendering techniques have achieved remarkable photorealism in 3D reconstruction mildenhall2020nerf ; kerbl2022efficient , their outputs often resemble “3D photographs”: they capture geometry and texture but lack true interactivity and structure. A truly effective digital replica must go beyond visual fidelity to offer semantically rich, simulation-ready, interactable environments. Such representations empower users to interact with, manipulate, and simulate complex scenes—capabilities essential for robotics, virtual and augmented reality, and embodied AI.
Compared to traditional 3D reconstruction techniques—such as Structure-from-Motion (SfM) schonberger2016structure ; wu2011visualsfm , Gaussian Splatting kerbl2022efficient , and Neural Radiance Fields (NeRF) mildenhall2020nerf ; barron2021mipnerf —we introduce the concept of graphics-ready reconstruction. This form of reconstruction goes beyond visual fidelity by incorporating additional structural and functional information to ensure compatibility with graphics pipelines and support downstream tasks such as interactive simulation. To achieve this, several key criteria must be satisfied. First, the scene should be object-centric, treating each object as an individual entity to enable physically plausible placement and manipulation. Second, object functionality must be modelled, allowing for interactive behaviors such as opening doors, drawers, or appliances. Third, the scene should use PBR materials to ensure photorealistic appearance under varying lighting and relighting conditions. Lastly, the integration of physical properties—including mass, gravity, and collision dynamics—is essential to support realistic interactions grounded in real-world physics. By meeting these requirements, we move beyond passive, geometry-focused reconstructions toward truly graphics-ready environments that combine visual realism with functional structure.
@@
To this end, we introduce LiteReality, a method that automatically transforms RGB-D scans into realistic, interactive indoor environments (see Figure 1). LiteReality is designed to handle the diverse challenges of real-life scans. At its core, LiteReality consists of four main stages: 1. Scene Perception and Parsing: Objects and room layouts are extracted from RGB-D scans using off-the-shelf methods and represented in a compact scene graph. This graph enforces spatial constraints and refines noisy input into a structured, physically plausible layout. 2. Object Reconstruction: Artist-designed 3D models that best match real-world objects are retrieved from a curated asset database. Though not perfectly aligned geometrically, these models provide high-fidelity geometry and support interactive functionality, including articulated components. 3. Material Painting: A robust material estimation and optimization framework assigns high-quality PBR materials to the retrieved 3D models, based on multiview RGB images. 4. Procedural Reconstruction: The final scene is assembled by constructing room structures, assembling doors and chairs, placing objects, and assigning basic physical properties and interaction logic—resulting in a lightweight, graphics-ready environment. We evaluate LiteReality on the ScanNet dataset dai2017scannet and on real-world indoor scans captured with an iPhone. Across diverse environments, our method consistently converts noisy, incomplete scans into compact, and realistic digital replicas. Our main contributions are as follows:
  • LiteReality framework: To the best of our knowledge, LiteReality is the first system that converts room-level scans into graphics-ready environments with both high realism and interactive functionality. We demonstrate good generalization across diverse real-world scans.
@@
  • Robust material painting: We introduce a novel material-painting approach that reliably transfers high-fidelity PBR materials onto 3D models—even in the presence of shape misalignments, challenging lighting, and diverse image styles. Quantitative evaluation shows SOTA performance.
@@ 2 Related Work
Figure 2: Pipeline of LiteReality. With input RGB-D scans, the process begins with scene perception and parsing, where room layouts and 3D object bounding box are detected and parsed into a structured and physically plausible arrangement with the help of a scene graph. In the object reconstruction stage, identical clustering first finds identical objects, and a hierarchical retrieval approach is conducted to match 3D models retrieved from the LiteReality database. The material painting stage retrieves and optimizes PBR materials by referencing observed images. Finally, procedural reconstruction assembles all the information into a graphics-ready environment that features realism and interactivity.
@@
While industries such as gaming, AR/VR, film production, and physical simulation rely heavily on traditional computer graphics pipelines to create 3D environments, most state-of-the-art 3D reconstruction methods mildenhall2020nerf ; schonberger2016structure ; kerbl2022efficient are not readily compatible with these pipelines. This incompatibility poses a barrier to their practical adoption in real-world applications. In the research domain, simulation environments designed for embodied AI training are the closest counterparts to graphics-ready environments, as they emphasize both interactivity and realism. Significant efforts have been made to develop such simulation platforms. AI2-THOR ai2thor , RoboTHOR robothor , OmniGibson li2023behavior , Habitat 3.0 puig2023habitat3 , ThreeDWorld threedworld , and VirtualHome puig2018virtualhome are notable examples. Despite their capabilities, these platforms often rely on manually crafted scenes, which limits scalability. Several works have proposed scalable scene generation methods. ProcTHOR procthor extends AI2-THOR with procedurally generated environments; Holodeck Yang_2024_CVPR creates scenes from natural language inputs; Phone2Proc phone2proc generates room layouts by placing semantically similar objects; and Digital Cousin dai2024acdc produces diverse simulated environments from single images. While these approaches improve scalability and diversity, they primarily focus on synthesizing virtual scenes for robotic training rather than faithfully reconstructing real-world environments. This highlights a key gap: building graphics-ready scenes that combine realistic appearance with physical interactivity. A concurrent work, Metascenes yu2025metascenesautomatedreplicacreation , is also working in a similar direction by proposing a large-scale dataset that contains a large number of simulatable scenes constructed from richly annotated ScanNet scans for embodied AI benchmarks. LiteReality, by contrast, is presented as a general-purpose tool for converting real-life scans to graphics-ready scenes—recovering artist-quality PBR materials for realism and generating fully interactive, realistic environments.
@@ 3.1 Scene Perception and Parsing
In the scene graph, nodes represent key elements in the scene, such as walls, windows, doors, and detected objects. Each node is associated with several attributes: 1) spatial properties, including center location $\mathbf{C}\in\mathbb{R}^{3}$ (where $\mathbf{C}=(x,y,z)$), dimensions $\mathbf{D}\in\mathbb{R}^{3}$ (where $\mathbf{D}=(w,h,l)$), and orientation $\boldsymbol{\theta}\in\mathbb{R}^{3}$ (three rotation angles around the $x$, $y$, and $z$ axes); and 2) appearance properties, represented by the top-$k$ most visible cropped images of the object, denoted as a set =$\mathcal{I}=\{I_{1},I_{2},\ldots,I_{k}\}$. These attributes are essential for material estimation, model selection, and layout optimization. Edges between nodes define key spatial relationships and, once constructed, can impose constraints during the spatial arrangement process, ensuring a physically plausible scene. We define four key spatial relationships: 1) attached to walls, for objects physically connected to walls (e.g., shelves, cabinets); 2) on top of, for objects placed on top of others (e.g., a vase on a table); 3) table-chair pair, grouping tables and chairs in typical arrangements; and 4) connecting to, where objects are next to each other and touched. These relationships provide essential constraints for maintaining consistent object placement, preventing collisions, and improving overall scene realism. To compute the edges of the scene graph, the construction process begins by parsing walls into connected components, forming closed polygonal regions that delineate individual rooms. Noisy wall detections are aligned by snapping them to a grid, ensuring well-defined room boundaries. Spatial relationships between objects are inferred based on their center location and orientation, with predefined rules governing each type of relationship. We resolve object collisions iteratively using a Constraint-Based Collision Resolution approach. For each intersecting pair, virtual forces are applied along collision vectors to separate them, while adhering to spatial constraints from the scene graph. For example, wall-attached objects are restricted to wall-aligned movement, and table-chair pairs retain their relative positions. This process continues until collisions are resolved or a maximum iteration count is reached, ensuring a realistic, physically plausible layout. More details can be found in the supplementary materials.
@@ 3.3 Material Painting Stage
3.4 Procedural Reconstruction Stage
Finally, we use a predefined procedural pipeline to reconstruct the scene. This begins with constructing the walls, followed by assembling windows and doors, and finally placing objects in the room. To support physics-based interaction, we assign rigid-body properties to each object, and use their mesh geometry as collision boundaries. Large structural elements, such as walls, are marked as passive rigid bodies, while movable items are set as active rigid bodies. Estimating object mass is crucial for realistic simulation. To achieve this, we crop images of each object and input them into a Multimodal Large Language Model (MLLM), which predicts their mass. These estimates are used to simulate realistic rigid-body dynamics. This entire procedural reconstruction process is conducted within Blender.
@@ Results.
Table 2: Object-Centric PBR Material Estimation Comparison. Evaluated on 110 objects across five real-world scanned scenes, LiteReality method demonstrates strong capabilities.
@@ 5 Conclusion
LiteReality demonstrates the ability to reconstruct realistic, interactive indoor scenes from RGB-D scans. Our results showcase the pipeline’s potential in generating visually compelling scenes with physics-based interactions and dynamic lighting. Future work will explore support for additional input modalities, such as single 360-degree images and egocentric videos. Another important direction is deeper physics integration and modelling for dynamic scene interactions and object augmentation. We hope this work serves as a strong foundation for graphics-ready 3D scene reconstruction, with broad applications in robotics, gaming, and AR/VR.
@@ C.1 Retrieval Similarity Evaluation Protocol
To compute Chamfer Distance, each CAD model is normalized—centered at the origin and scaled to fit within a unit cube (1$\times$1$\times$1). We uniformly sample 10,000 surface points from both the retrieved and ground-truth CAD models. The L1 Chamfer Distance is calculated bidirectionally as follows:
@@ Appendix E More Results
(a) Chair, Fireplace, Sofa, Storage
@@
(b) Table, Television, Door, Windows
@@
Table 5: Per-category evaluation metrics (RMSE ↓, SSIM ↑, LPIPS ↓) for different methods across various object types.
@@ Appendix F Application of LiteReality
Appendix G Limitations and Future Work
Limitations
While LiteReality converts RGB–D scans into compact, interactive 3D scenes, it still has several limitations. First, our procedural reconstruction relies on hand-tuned rules and artist-crafted assets, which can fail in unconventional or highly customized layouts (e.g., a sink flush-mounted within bespoke cabinetry). Second, although our Material Painting module recovers high-quality, spatially varying albedo maps, it does not estimate illumination or specular properties, limiting relighting fidelity under novel lighting conditions. Third, small or cluttered objects—such as mugs, books, and cables—are currently omitted, reducing scene completeness. Fourth, reconstruction quality remains tightly coupled to upstream perception: errors in object detection can propagate. Finally, our simple constraint-based collision solver can struggle in complex arrangements, occasionally producing interpenetrations and preventing fully physics-valid simulations.
Future Work
Integrating small-object detection and retrieval into the pipeline would enrich scene detail and realism. Incorporating learning-based light estimation together with full SVBRDF recovery would enable faithful relighting under novel illuminations. Exploring learned or physics-driven collision-resolution techniques could ensure non-overlapping, simulatable layouts. Finally, scaling the approach from individual rooms to multi-room and building-scale reconstructions would unlock applications in large-scale visualization and digital-twin creation.

### v2

# arXiv:2507.02861v2[cs.CV] 27 Feb 2026

@@ Abstract
We propose LiteReality, a novel pipeline that converts RGB-D scans of indoor environments into compact, realistic, and interactive 3D virtual replicas. LiteReality not only reconstructs scenes that visually resemble reality but also supports key features essential for graphics pipelines—such as object individuality, articulation, high-quality physically based rendering materials. At its core, LiteReality first performs scene understanding and parses the results into a coherent 3D layout and objects, with the help of a structured scene graph. It then reconstructs the scene by retrieving the most visually similar 3D artist-crafted models from a curated asset database. Later, the Material Painting module enhances the realism of retrieved objects by recovering high-quality, spatially varying materials. Finally, the reconstructed scene is integrated into a simulation engine with basic physical properties applied to enable interactive behavior. The resulting scenes are compact, editable, and fully compatible with standard graphics pipelines, making them suitable for applications in AR/VR, gaming, robotics, and digital twins. In addition, LiteReality introduces a training-free object retrieval module that achieves state-of-the-art similarity performance, as benchmarked on the Scan2CAD dataset, along with a robust Material Painting module capable of transferring appearances from images of any style to 3D assets—even in the presence of severe misalignment, occlusion, and poor lighting. We demonstrate the effectiveness of LiteReality on both real-life scans and public datasets.
Figure 1: From RGB-D Scan to Graphics-Ready 3D Scene. LiteReality reconstructs compact, realistic 3D environments from real-world RGB-D scans, featuring near-photorealistic appearance, articulated geometry, and physically based rendering (PBR) materials—providing assets that can be easily integrated into simulation or rendering pipeline.
@@ 1 Introduction
Creating digital replicas of real-world environments remains a central challenge in machine perception, computer vision, and graphics. Although differentiable rendering techniques have achieved remarkable photorealism in 3D reconstruction mildenhall2020nerf ; kerbl2022efficient , their outputs often resemble “3D photographs”: they capture geometry and texture but lack true interactivity and structure. A truly effective digital replica must go beyond visual fidelity to offer semantically rich, simulation-ready, interactive environments. Such representations empower users to interact with, manipulate, and simulate complex scenes—capabilities essential for robotics, virtual and augmented reality, and embodied AI.
Compared to traditional 3D reconstruction techniques—such as Structure-from-Motion (SfM) schonberger2016structure ; wu2011visualsfm , Gaussian Splatting kerbl2022efficient , and Neural Radiance Fields (NeRF) mildenhall2020nerf ; barron2021mipnerf —we introduce the concept of graphics-ready reconstruction. This form of reconstruction goes beyond visual fidelity by incorporating additional structural and functional information to ensure compatibility with graphics pipelines and support downstream tasks such as interactive simulation. To achieve this, several key criteria must be satisfied. First, the scene should be object-centric, treating each object as an individual entity to enable physically plausible placement and manipulation. Second, object functionality must be modeled, allowing for interactive behaviors such as opening doors, drawers, or appliances. Third, the scene should use PBR materials to ensure photorealistic appearance under varying lighting conditions. Lastly, the integration of physical properties—including mass, gravity, and collision dynamics—is essential to support realistic interactions grounded in real-world physics. By meeting these requirements, we move beyond passive, geometry-focused reconstructions toward truly graphics-ready environments that combine visual realism with functional structure.
@@
To this end, we introduce LiteReality, a method that automatically transforms RGB-D scans into realistic, interactive indoor environments (see Figure 1). LiteReality is designed to handle the diverse challenges of real-life scans. At its core, LiteReality consists of four main stages: 1. Scene Perception and Parsing: Objects and room layouts are extracted from RGB-D scans using off-the-shelf methods apple_roomplan_2022 ; lazarow2024cubify and represented in a compact scene graph. This graph enforces spatial constraints and refines noisy input into a structured, physically plausible layout. 2. Object Reconstruction: Artist-designed 3D models that best match real-world objects are retrieved from a curated asset database. Though not perfectly aligned geometrically, these models provide high-fidelity geometry and support interactive functionality, including articulated components. 3. Material Painting: A robust material estimation and optimization framework assigns high-quality PBR materials to the retrieved 3D models, based on multi-view RGB images. 4. Procedural Reconstruction: The final scene is assembled by constructing room structures, assembling doors and chairs, placing objects, and assigning basic physical properties and interaction logic—resulting in a lightweight, graphics-ready environment. We evaluate LiteReality on the ScanNet dataset dai2017scannet and on real-world indoor scans captured with an iPhone. Across diverse environments, our method consistently converts noisy, incomplete scans into compact, and realistic digital replicas. Our main contributions are as follows:
  • LiteReality framework: To the best of our knowledge, LiteReality is the first system capable of reconstructing room-level RGB-D scans into compact and realistic CAD representations that feature high-quality PBR materials, providing a strong foundation for downstream applications such as rendering, simulation, and virtual interaction. The system generalizes well across a wide range of real-world scenes.
@@
  • Robust material painting: We introduce a novel material-painting approach that reliably transfers high-fidelity PBR materials onto 3D models—even in the presence of shape misalignment, challenging lighting, and diverse image styles. Quantitative evaluation shows SOTA performance.
@@ 2 Related Work
Figure 2: Pipeline of LiteReality. Given input RGB-D scans, the process begins with scene perception and parsing, where room layouts and 3D object bounding boxes are detected and organized into a structured, physically plausible arrangement using a scene graph. In the object reconstruction stage, identical clustering first identifies repeated objects, followed by a hierarchical retrieval procedure that matches 3D models from the LiteReality database. The material painting stage retrieves and optimizes PBR materials by referencing the observed images. Finally, the procedural reconstruction stage assembles all components into a graphics-ready environment featuring realistic appearance and seamless integration with standard graphics pipelines.
@@
While industries such as gaming, AR/VR, film production, and physical simulation rely heavily on traditional computer graphics pipelines to create 3D environments, most state-of-the-art 3D reconstruction methods mildenhall2020nerf ; schonberger2016structure ; kerbl2022efficient are not readily compatible with these pipelines. This incompatibility poses a barrier to their practical adoption in real-world applications. In the research domain, simulation environments designed for embodied AI training are the closest counterparts to graphics-ready environments, as they emphasize both interactivity and realism. Significant efforts have been made to develop such simulation platforms. AI2-THOR ai2thor , RoboTHOR robothor , OmniGibson li2023behavior , Habitat 3.0 puig2023habitat3 , ThreeDWorld threedworld , and VirtualHome puig2018virtualhome are notable examples. Despite their capabilities, these platforms often rely on manually crafted scenes, which limits scalability. Several works have proposed scalable scene generation methods. ProcTHOR procthor extends AI2-THOR with procedurally generated environments; Holodeck Yang_2024_CVPR creates scenes from natural language inputs; Phone2Proc phone2proc , generates room layouts by placing semantically similar objects; and Digital Cousin dai2024acdc produces diverse simulated environments from single images. More recent contributions in this area include Scenethesis ling2024scenethesis , PAT3D lin2025pat3dphysicsaugmentedtextto3dscene and VIGA yin2026visionasinversegraphicsagentinterleavedmultimodal . While these approaches improve scalability and diversity, they primarily focus on synthesizing virtual scenes for robotic training rather than faithfully reconstructing real-world environments. This highlights a key gap: building graphics-ready scenes that combine realistic appearance with physical interactivity. A concurrent work, Metascenes yu2025metascenesautomatedreplicacreation , is also working in a similar direction by proposing a large-scale dataset that contains a large number of simulatable scenes constructed from richly annotated ScanNet scans for embodied AI benchmarks. LiteReality, by contrast, is presented as a general-purpose tool for converting real-life scans to graphics-ready scenes.
@@ 3.1 Scene Perception and Parsing
In the scene graph, nodes represent key elements in the scene, such as walls, windows, doors, and detected objects. Each node is associated with several attributes: 1) spatial properties, including center location $\mathbf{C}\in\mathbb{R}^{3}$ (where $\mathbf{C}=(x,y,z)$), dimensions $\mathbf{D}\in\mathbb{R}^{3}$ (where $\mathbf{D}=(w,h,l)$), and orientation $\boldsymbol{\theta}\in\mathbb{R}^{3}$ (three rotation angles around the $x$, $y$, and $z$ axes); and 2) appearance properties, represented by the top-$k$ most visible cropped images of the object, denoted as a set $\mathcal{I}=\{I_{1},I_{2},\ldots,I_{k}\}$. These attributes are essential for material estimation, model selection, and layout optimization. Edges between nodes define key spatial relationships and, once constructed, can impose constraints during the spatial arrangement process, ensuring a physically plausible scene. We define four key spatial relationships: 1) attached to walls, for objects physically connected to walls (e.g., shelves, cabinets); 2) on top of, for objects placed on top of others (e.g., a vase on a table); 3) table-chair pair, grouping tables and chairs in typical arrangements; and 4) connecting to, where objects are next to each other and touched. These relationships provide essential constraints for maintaining consistent object placement, preventing collisions, and improving overall scene realism. To compute the edges of the scene graph, the construction process begins by parsing walls into connected components, forming closed polygonal regions that delineate individual rooms. Noisy wall detections are aligned by snapping them to a grid, ensuring well-defined room boundaries. Spatial relationships between objects are inferred based on their center location and orientation, with predefined rules governing each type of relationship. We resolve object collisions iteratively using a Constraint-Based Collision Resolution approach. For each intersecting pair, virtual forces are applied along collision vectors to separate them, while adhering to spatial constraints from the scene graph. For example, wall-attached objects are restricted to wall-aligned movement, and table-chair pairs retain their relative positions. This process continues until collisions are resolved or a maximum iteration count is reached, ensuring a realistic, physically plausible layout. More details can be found in the supplementary materials.
@@ 3.3 Material Painting Stage
3.4 Procedural Reconstruction Stage in Blender
Finally, we use a predefined procedural pipeline to reconstruct the scene. The process begins with constructing the walls, followed by assembling windows and doors, and finally placing objects within the room. To support physics-based interaction, we assign rigid-body properties to each object and use their mesh geometry as collision boundaries. Large structural elements, such as walls, are marked as passive rigid bodies, while movable items are set as active rigid bodies. Estimating object mass is crucial for realistic simulation. To achieve this, we crop images of each object and input them into a Multimodal Large Language Model (MLLM), which predicts their mass—later used to simulate realistic rigid-body dynamics. The entire procedural reconstruction process is carried out within Blender.
@@ Results.
Table 2: Object‐Centric PBR Material Estimation Comparison. Evaluated on 110 objects across five real‐world scanned scenes, LiteReality method demonstrates strong capabilities.
@@
Runtime Analysis
On a single NVIDIA RTX 3090 GPU with 24 GB of memory, the complete reconstruction of a room-scale scene takes between 20 and 60 minutes, depending on scene complexity. For example, a small study or bedroom scene containing around 10–15 objects typically requires about 20 minutes, whereas larger spaces such as meeting rooms or boardrooms with 40–50 objects take up to an hour. Among the individual stages, preprocessing and scene parsing complete within 1–3 minutes, object retrieval requires 2–5 minutes, while the material painting stage dominates the runtime, taking roughly 15–50 minutes depending on the number of objects. Procedural reconstruction and export in Blender add less than two minutes on average.
@@ 5 Conclusion
LiteReality demonstrates the ability to reconstruct realistic and interactive indoor scenes from RGB-D scans. Our results highlight the potential of the proposed pipeline in generating visually compelling environments with physics-based interactions and dynamic lighting. We hope this work will inspire further exploration toward graphics-ready 3D scene reconstruction.
Limitations.
The overall reconstruction quality remains closely tied to the accuracy of upstream perception. Errors in object detection and layout estimation can propagate to later stages. In most real-world cases, our input scans are captured using Apple’s RoomPlan app, which provides reasonably reliable geometry through interactive scanning. While we propose a comprehensive retrieval pipeline, object retrieval remains inherently challenging—an occasional mismatch can noticeably affect scene realism. Fortunately, due to the object-centric nature of our representation, such errors are easy to correct through local replacement or refinement. Finally, our current constraint-based collision solver may struggle in densely packed or complex arrangements, sometimes leading to interpenetrations and preventing fully physics-valid simulations.
Future Work.
Future work could focus on extending object detection and retrieval to include smaller items, thereby enriching scene detail and overall realism. Incorporating learning-based light estimation together with full SVBRDF recovery would enable faithful relighting under novel illumination conditions. Scaling the pipeline from single-room to multi-room and building-level reconstructions would further unlock applications in large-scale visualization, robotics simulation, and digital-twin creation. In addition, optimizing the entire system for real-time or near real-time performance would make it highly practical for interactive applications and on-device deployment.
Acknowledgment.
This work was supported by the Girton College Graduate Research Award at the University of Cambridge, the School of Technology, and the EPSRC Centre for Doctoral Training. We gratefully acknowledge the insightful discussions and feedback from Prof. Kwang Moo Yi, Erqun Dong, Weiyang Liu, Prof. Ayush Tewari, and Shangzhan Zhang throughout the course of this work. We also appreciate the kind support of Prof. Angela Dai for providing the voiceover for the demo videos, and we thank Jieying Chen for her help in proofreading the paper.
@@ C.1 Retrieval Similarity Evaluation Protocol
To compute Chamfer Distance, each CAD model is normalized—centered at the origin and scaled to fit within a unit cube ($\times1$$\times$1). We uniformly sample 10,000 surface points from both the retrieved and ground-truth CAD models. The L1 Chamfer Distance is calculated bidirectionally as follows:
@@ Appendix E More Results
(a) Chair, Fireplace, Sofa, and Storage
@@
(b) Table, Television, Door, and Windows
@@
Table 5: Per‐category evaluation metrics (RMSE ↓, SSIM ↑, LPIPS ↓) for different methods across representative object types. Best results are in bold.